The second side of a triangular deck is 4 feet longer than the shortest side and a third side that is 4 feet shorter than twice the length of the shortest side. If the perimeter of the deck is 48 ​feet, what are the lengths of the three​ sides?
Thank you SO much for the help!

Answer by ankor@dixie-net.com(22740)

let s = the length of the shortest side
:
Write an equation for each statement
:
The second side of a triangular deck is 4 feet longer than the shortest side
(s+4) = the 2nd side
and a third side that is 4 feet shorter than twice the length of the shortest side.
(2s-4) = the 3rd side
:
If the perimeter of the deck is 48 ​feet, what are the lengths of the three​ sides?
s + (s+4) + (2s-4) = 48
Combine like terms
s + s + 2s + 4 - 4 = 48
4s = 48
s = 48/4
s = 12 ft is the shortest side
:
:
I'll let you find the 2nd and 3rd sides, ensure they add up to 48